40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 5 07:19:16 2012, kiwamu, Update, LSC, DRMI locking DRMI_2012Apr4_edit.png
    Reply  Sat Apr 7 01:56:05 2012, kiwamu, Update, LSC, OSA signal in DRMI condition OSA.png
       Reply  Sat Apr 7 06:58:34 2012, kiwamu, Update, LSC, DRMI lock : lost good alignment 
Message ID: 6489     Entry time: Thu Apr 5 07:19:16 2012     Reply to this: 6506
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: DRMI locking 

 I tried locking the DRMI to the signal-extraction condition with the new trigger by AS110.

A first thing I tried was : flipping the control sign of the SRCL while keeping the same control setups for the PRCL and MICH.

Occasionally the DRMI was "sort of" locked and hence I believe this setup must be a good starting point.

As a next step I will try some different gains and demodulation phase to make it more lockable.

 


(Time series)

DRMI_2012Apr4_edit.png

 The picture above is time series of some signals when the DRMI was barely locked.
The red arrows indicate the durations when the DRMI was sort of locked.
 (Green curve) REFLDC becoming a high value state, which indicates that the carrier is anti-resonant.
 (Red curve) ASDC becoming dark, which indicates the MICH is in the vicinity of the dark condition.
 (Brown curve) AS110 becoming a high value state, which means the 55 MHz sidebands got amplified by the SRCL.
 (Blue curve) POP22 becoming a high value state, which indicates that the 11 MHz sidebands are resonating in the PRC.
 
According to the measurement of AS110 when PRMI was locked (#6488), the AS110 signal went up to ~ 1 counts or so.
On the other hand when the DRMI was locked the AS110 went to up more than 10 counts as shown in the plot above.
Therefore at least some kind of signal amplification is happening for the 55 MHz sidebands in the SRC.
Looking at the AS CCD, I found that the beam looked like a TEM01 mode (two beam spots at top and bottom) every time when the DRMI was locked.
 
(settings)
  • REFL33I => PRCL  G = -0.2
  • AS55Q => MICH    G = -6
  • AS55I => SRCL     G = 1   (G = -50 for the signal recycling condition)
  • AS55 demod phase = 17 deg
ELOG V3.1.3-